OpenStudy (anonymous):

What is the product of 2the square root of the quantity 15 times k. • 6kthe square root of 3.? Simplify if possible.

OpenStudy (anonymous):

\[2\sqrt{15k} \times 6k \sqrt{3}\]

OpenStudy (anonymous):

hint: \[a \sqrt{b}.c \sqrt{d}=ac \sqrt{bd}\]

OpenStudy (anonymous):

so? 12k sqrt 45k? that doesnt make sense..:p

OpenStudy (anonymous):

why not,then try if you can simplify the square root further

OpenStudy (anonymous):

but the anwsers give me choices like 36 and 60 start with?

OpenStudy (anonymous):

yes that's why you have to simplify further another hint 45=9.5

OpenStudy (anonymous):

um what? lol

OpenStudy (anonymous):